The nine employes of the Ellington Dance Studio each earn $600 per week. A new employee is hired who earns $1500 per week.
Describe how the new employee's earnings will affect the mean and median earnings of Ellington Studio's employees. Use mathematics to support your answer
1
Expert's answer
2012-03-27T11:02:54-0400
Ex- mean& for nine employes =600- 'cause all receive so. Median is found& by the formula (600+600)/2 = 600 as well.
Ex afterwise is 600*9/10+1500*1/10=(5400+1500)/10=690- the new mean
9/10 –probability to choose person with salary 600
1/10 –probability to choose person with salary 1500
Median will be the same- by formula (600+600)/2